OpenStudy (lgbasallote):

DIFFERENTIAL EQUATIONS QUESTION: Evaluate L{t^2 + 4t - 5} can i do it like L{t^2} + L{4t} - L{5}?

OpenStudy (anonymous):

I guess the L here denotes the La Place Transform? \[ \Large \mathcal{L}\lbrace t^2+4t-5\rbrace \]

OpenStudy (lgbasallote):

yup

OpenStudy (anonymous):

The La Place Transform is a linear operator. \[ \Large \mathcal{L}\lbrace c_1f(t)+c_2g(t)\rbrace= \int_0^\infty e^{-st}(c_1f(t)+c_sg(t))dt \\= \Large c_1 \int_'^\infty e^{-st}f(t)+c_2\int_0^\infty e^{-st}g(t)dt \\ =\Large c_1 \mathcal{L}\lbrace f(t)\rbrace + c_2 \mathcal{L} \lbrace g(t)\rbrace \]

OpenStudy (anonymous):

therefore, yes you can.

OpenStudy (lgbasallote):

o.O

OpenStudy (lgbasallote):

...so i can right?

OpenStudy (lgbasallote):

L{t^2} = 2/s^3 right?

OpenStudy (anonymous):

yes, it's a linear operator and right, if I remember it correctly, I usually use a cheat sheet *laughs*

OpenStudy (anonymous):

Let me check.

OpenStudy (lgbasallote):

well i use \[L{t^n} \implies \frac{n!}{s^{n+1}}\] lol

OpenStudy (anonymous):

\[ \Large t^n = \frac{n!}{s^{n+1}} \]

OpenStudy (anonymous):

yes, you're correct.

OpenStudy (anonymous):

http://tutorial.math.lamar.edu/pdf/Laplace_Table.pdf I should save that somewhere.

OpenStudy (lgbasallote):

and L{4t} would be 4\s^2 right?

OpenStudy (anonymous):

correct, you can pull the 4 out of the laplace transform and leave it as a scalar identity.

OpenStudy (lgbasallote):

so \[\huge L\left[t^2 + 4n - 5 \right] \implies \frac{2}{s^3} + \frac{4}{s^2} - \frac 5s\] right? idk how to do those curly brackets in latex lol

OpenStudy (anonymous):

Curly brackets are .\ lbrace and .\ rbrace, the Laplace Transform is written as .\ mathcal{L} if you want to be rigorous about it

OpenStudy (anonymous):

And yes, you got it.

OpenStudy (lgbasallote):

sweet

OpenStudy (anonymous):

(-;